OpenStudy (anonymous):

The function f(x) is shown below. f(x) = (2)x+1 The function f(x) is shifted to the left by 13 units. Which of the following best represents the new function? f(x) = (2)x + 13 f(x) = (2)x + 14 f(x) = (2)x -12 f(x) = (2)x - 13

jimthompson5910 (jim_thompson5910):

hint: to shift f(x) to the left by 13 units, you have to replace x with x-13. So you need to evaluate f(x-13)

OpenStudy (anonymous):

thank you!! :) I hav another question if u could help me with

OpenStudy (anonymous):

Wendy enjoys listening to hip hop and downloads music regularly from the internet. She already has four hours of hip hop in her collection. From the second month onward, the number of hours of hip hop music she downloads each month is nine percent greater than the number of hours of hip hop music she had downloaded the previous month. Which function best shows the total number of hours of hip hop music Wendy has in her collection f(x) after x months? f(x) = (1.09)x + 3 f(x) = 3(1.09)x f(x) = (1.09)x + 2 f(x) = (1.09)x - 2

jimthompson5910 (jim_thompson5910):

Odd...it mentions '4' but I don't see it anywhere in any of the functions.

jimthompson5910 (jim_thompson5910):

You sure it's not "3 hours" instead of "4 hours"?

OpenStudy (anonymous):

idk it says four

OpenStudy (anonymous):

can u help me with this one instead Which exponential function goes through the points (1, 8) and (4, 64)? f(x) = 4(2)x f(x) = 2(4)x f(x) = 4(2)-x f(x) = 2(4)-x

jimthompson5910 (jim_thompson5910):

plug x = 1 into each function and evaluate if you get y = 8, then you're on the right track if you don't get y = 8, then you can eliminate that answer choice

jimthompson5910 (jim_thompson5910):

same thing for x = 4 you should get y = 64 if you don't, then eliminate that choice

OpenStudy (anonymous):

is it f(x) = 2(4)x

jimthompson5910 (jim_thompson5910):

what happens when x = 4?

OpenStudy (anonymous):

it not that one because it becomes 32 right

jimthompson5910 (jim_thompson5910):

the function is f(x) = 2(4)^x

jimthompson5910 (jim_thompson5910):

not f(x) = 2(4)x

OpenStudy (anonymous):

i don't understand this

jimthompson5910 (jim_thompson5910):

f(x) = 2(4)^x f(4) = 2(4)^4 f(4) = ???

OpenStudy (anonymous):

2*4=8 so 8*8*8*8?

jimthompson5910 (jim_thompson5910):

what is 4^4?

OpenStudy (anonymous):

256

jimthompson5910 (jim_thompson5910):

now multiply that by 2

OpenStudy (anonymous):

512

jimthompson5910 (jim_thompson5910):

so f(4) = 512, but we wanted f(4) to be equal to 64

OpenStudy (anonymous):

so what do we need to do now

jimthompson5910 (jim_thompson5910):

so clearly (4,64) is NOT on the graph of f(x) = 2(4)^x

OpenStudy (anonymous):

so it's the first one because f(4) = 4(2)4= 4*4=16 16*4=64 and f(1) = 4(2)1=1*2=2 and 2*4=8

jimthompson5910 (jim_thompson5910):

it's a bit hard to follow your steps, but yeah it's the first one

OpenStudy (anonymous):

Thank you so much i just have one more question The table below shows the values of f(n) for different values of n. n 1 2 3 4 5 6 f(n) 1 2 2 4 8 32 Which recursive function best represents the values shown in the table? f(1) = 1, f(2) = 2, f(n) = f(n -1) f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = f(n -1) + f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = f(n -3) + f(n - 2); n > 2 f(1) = 1, f(2) = 2, f(n) = f(n -3) f(n - 2); n > 2

jimthompson5910 (jim_thompson5910):

do you know how recursive functions work?

OpenStudy (anonymous):

not that well

jimthompson5910 (jim_thompson5910):

what is f(3) in this case (look at the table)

OpenStudy (anonymous):

idk

jimthompson5910 (jim_thompson5910):

look for n = 3, then look below that

OpenStudy (anonymous):

the 2

jimthompson5910 (jim_thompson5910):

so f(3) = 2

jimthompson5910 (jim_thompson5910):

so let's say the function was f(n) = f(n -1) + f(n - 2) if that were true, then plugging in n = 3 gives us f(n) = f(n -1) + f(n - 2) f(3) = f(3 -1) + f(3 - 2) ... plug in n = 3 (ie replace each n with 3) f(3) = f(2) + f(1) what's the next step?

OpenStudy (anonymous):

u add f(2) and f(1)

jimthompson5910 (jim_thompson5910):

what is f(2) and f(1)?

jimthompson5910 (jim_thompson5910):

they are given in the table

OpenStudy (anonymous):

1 and 2

OpenStudy (anonymous):

idk this stuff

jimthompson5910 (jim_thompson5910):

that means f(3) = f(2) + f(1) becomes f(3) = 2+1 f(3) = 3 but remember we wanted f(3) to be equal to 2 (and not 3)

OpenStudy (anonymous):

so it needs to be subtraction

jimthompson5910 (jim_thompson5910):

so that means that f(n) = f(n -1) + f(n - 2) cannot be the answer why not? because when you plug in n = 3 you get f(3) = 3 when you should get f(3) = 2

jimthompson5910 (jim_thompson5910):

see how I eliminated that answer choice?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

f(x)=2x - 12 because you shift to the left 13=> 1-13=-12

OpenStudy (anonymous):

What's the answer?